OpenStudy (anonymous):

help me will GIVE MEDALS At the beginning of the summer, Jenna had saved $105. To earn money over the summer, she is working part time at the local community center. She will earn $35 per day of work. If x represents the number of days she works, which of the following equations represents how much money she will have, including her savings? A.y = 105 - 35x B.y = x + 105 C.y = 105x + 35 D.y = 35x + 105

OpenStudy (rccb):

D) y=35x+105

OpenStudy (anonymous):

thank you can you help me with some more

OpenStudy (rccb):

sure

OpenStudy (anonymous):

On Fridays at the ballpark, a cup of soda costs $4, and each refill costs $0.60. If x represents the number of refills, which of the following equations represents the amount a person spends on soda? A.y = 0.6x + 4 B.y = 4x + 0.6 C.y = 0.6x + 4.6 D.y = 4x + 4.6

OpenStudy (rccb):

which do you think it is?

OpenStudy (anonymous):

C

OpenStudy (rccb):

Close, the first time you buy the drink its the cup with soda in it, so that is 4 dollars to begin with, and every time you fill it up after that, it is 60 cents more(0.6)

OpenStudy (rccb):

the number of times you fill it up is x, so 0.6x would be part of the equation

OpenStudy (anonymous):

ok then its B right ?

OpenStudy (rccb):

no, x is the number of times you refill the cup, and they charge 60 cents for each refill, so if you refill it 5 times then it would be 0.6 times 5, but since we don't know how many times we refill the cup, the number stays 0.6x

OpenStudy (anonymous):

oh ok i get it

OpenStudy (anonymous):

its A

OpenStudy (rccb):

yes

OpenStudy (rccb):

you would take that number and add it to the first price(4) and that would equal y

OpenStudy (anonymous):

d

OpenStudy (rccb):

a, they only charge 4 dollars the first time, each refill is .60

OpenStudy (anonymous):

o k i inbox you

OpenStudy (anonymous):

i dont need help any more thank you @rccb